Mrs. Connor's physics class has 105 students, classified by academic year and gender, as ilustrated in the table. Mrs. Connor randomly chooses one student to collect yesterday's work. 

Mrs. Connor's Physics Class  

Academic YearMalesFemales
Freshmen1711
Sophomores135
Juniors915
Seniors1619

Step 1 of 2: What is the probability that she selects a female, given that she chooses randomly from only the freshmen Express your answer as a fraction or a decimal rounded to four decimal places.
Answer #1

No of freshman= 17+11 =28

No of females freshman=11

Hence probability of female getting kicked from freshmen=11/28= 0.3929